# Multi-Objective Coverage Bayesian Optimization (MOCOBO)

### v3

@@ 8. Answer: [Yes]
  9. Justification: See Section˜A.4 for discussion of limitations.
@@ 13. Answer: [Yes]
  14. Justification: All theoretical claims can be found in Section˜3.1.2. All sets of assumptions and full proofs of each claim are provided in Section˜3.1.2, Section˜A.9, and Section˜A.10. As far as we are aware, each proof is complete and correct.
@@ 18. Answer: [Yes]
  19. Justification: We provide detailed explanation of how our method works in Section˜3 and all additional required details to reproduce results in Section˜4 and Section˜A.7. Additionally, we provide a link to a public GitHub repository containing the source code used to run our method and produce results provided in this paper. This GitHub repository contains organized code that will allow any reader to run MOCOBO on all tasks in this paper. Additionally, the README in the repository provides detailed instructions to make setting up the proper environment and running the code easy for users.
@@ 28. Answer: [Yes]
  29. Justification: All chosen hyper-parameters and implementation details are stated in section 4 and Section˜A.7.
@@ 38. Answer: [Yes]
  39. Justification: All compute details are provided in Section˜A.6.
@@ 48. Answer: [Yes]
  49. Justification: See discussion of broader societal impacts in Section˜A.5.

### v4

@@ 8. Answer: [Yes]
  9. Justification: See Appendix˜D for discussion of limitations.
@@ 13. Answer: [Yes]
  14. Justification: All theoretical claims can be found in Section˜3.2.2. All sets of assumptions and full proofs of each claim are provided in Section˜3.2.2, Appendix˜I, and Appendix˜J. As far as we are aware, each proof is complete and correct.
@@ 18. Answer: [Yes]
  19. Justification: We provide detailed explanation of how our method works in Section˜3 and all additional required details to reproduce results in Section˜4 and Appendix˜G. Additionally, we provide a link to a public GitHub repository containing the source code used to run our method and produce results provided in this paper. This GitHub repository contains organized code that will allow any reader to run MOCOBO on all tasks in this paper. Additionally, the README in the repository provides detailed instructions to make setting up the proper environment and running the code easy for users.
@@ 28. Answer: [Yes]
  29. Justification: All chosen hyper-parameters and implementation details are stated in section 4 and Appendix˜G.
@@ 38. Answer: [Yes]
  39. Justification: All compute details are provided in Appendix˜F.
@@ 48. Answer: [Yes]
  49. Justification: See discussion of broader societal impacts in Appendix˜E.